compute the discriminant of the quadratic equation 2x^2-3x-5=0 and then write a brief sentence describing the number and type of solutions for the equation. show your work
Answer by jorel1380(3719)

2x^2-3x-5=0
(2x-5)(x+1)=0
x=5/2 or -1..